OpenStudy (anonymous):

Can someone please explain this Trig question, Find sec(x) in terms of sin(x)

jimthompson5910 (jim_thompson5910):

hints: \[\Large \sec(x) = \frac{1}{\cos(x)}\] \[\Large \sin^2(x) + \cos^2(x) = 1\]